OpenStudy (anonymous):

\[a_{k+2}=a_{k+1}\frac{k+1}{k+2}\] \[a_{1}\frac{1}{2} +a_{2}\frac{2}{3} +a_{3}\frac{3}{4} +a_{4}\frac{4}{5} +a_{5}\frac{5}{6} +a_{6}\frac{6}{7} +a_{7}\frac{7}{8} +a_{8}\frac{8}{9} +a_{9}\frac{9}{10} +a_{10}\frac{10}{11} +a_{11}\frac{11}{12}\]

zepdrix (zepdrix):

I'm not sure where you're going with the sum, you're jumping ahead too much and losing track of where we were :)

OpenStudy (anonymous):

oh and we said \[a_2=\frac{a_1}{2}\]

zepdrix (zepdrix):

Get your list straightened out first, everything in terms of A1.

OpenStudy (anonymous):

\[\frac{a_1}{2} +\frac{a_1}{3}+\dots\]

OpenStudy (anonymous):

I forgot how to do a_3

zepdrix (zepdrix):

\[k=1 \quad \rightarrow a_3=2a_2/3 \quad \rightarrow \quad a_3=a_1/3\]\[k=2 \quad \rightarrow a_4=3a_3/4 \quad \rightarrow \quad a_4=a_1/4\]\[k=3 \quad \rightarrow a_5=4a_4/5 \quad \rightarrow \quad a_5=a_1/5\]

zepdrix (zepdrix):

This is the list that we're generating.

zepdrix (zepdrix):

Confused how I got from the middle step to the last?

OpenStudy (anonymous):

Wow...I did get a little ahead of myself.... No it makes perfect sense

zepdrix (zepdrix):

You're forgetting all the X values in your series solution. Remember when we went over that before? :) Our solution should be of the form,\[\large \sum_{n=0}^\infty a_n x^n\]So when you expand that out, you should be getting powers of X along with your A's

OpenStudy (anonymous):

\[k=1 \quad \rightarrow a_3=2a_2/3 \quad \rightarrow \quad a_3=a_1/3\] \[k=2 \quad \rightarrow a_4=3a_3/4 \quad \rightarrow \quad a_4=a_1/4\] \[k=3 \quad \rightarrow a_5=4a_4/5 \quad \rightarrow \quad a_5=a_1/5\] what does \quad do?

zepdrix (zepdrix):

it makes a nice big space. \qquad makes a larger space. \; makes a small space.

OpenStudy (anonymous):

I see

zepdrix (zepdrix):

Oh actually, since we started with y' as lowest degree of y, i think our solution will start at n=1.... Otherwise we end up with an A0 term that we can't use.

OpenStudy (anonymous):

\[k=1 \quad \rightarrow a_3=2\frac{a_2}{3} \quad \rightarrow \quad a_3=\frac{a_1}{3}\] \[k=2 \quad \rightarrow a_4=3\frac{a_3}{4} \quad \rightarrow \quad a_4=\frac{a_1}{4}\] \[k=3 \quad \rightarrow a_5=4\frac{a_4}{5} \quad \rightarrow \quad a_5=\frac{a_1}{5}\] \[k=4 \quad \rightarrow a_6=5\frac{a_5}{6} \quad \rightarrow \quad a_6=\frac{a_1}{6}\] \[k=5 \quad \rightarrow a_7=6\frac{a_6}{7} \quad \rightarrow \quad a_7=\frac{a_1}{7}\] \[k=6 \quad \rightarrow a_8=7\frac{a_7}{8} \quad \rightarrow \quad a_8=\frac{a_1}{8}\] \[k=7 \quad \rightarrow a_9=8\frac{a_8}{9} \quad \rightarrow \quad a_9=\frac{a_1}{9}\] \[k=8 \quad \rightarrow a_{10}=9\frac{a_9}{10} \quad \rightarrow \quad a_{10}=\frac{a_1}{10}\] \[k=9 \quad \rightarrow a_{11}=10\frac{a_{11}}{12} \quad \rightarrow \quad a_{11}=\frac{a_1}{11}\] \[k=10 \quad \rightarrow a_{12}=12\frac{a_{12}}{13} \quad \rightarrow \quad a_{12}=\frac{a_1}{12}\]

zepdrix (zepdrix):

Yah I think you have the right pattern! That's what I was coming up with at least.

zepdrix (zepdrix):

\[y\quad =\quad \sum_{n=1}^\infty a_n x^n \quad =\]\[a_1x^1+a_2x^2+a_3x^3+a_4x^4+...\]

zepdrix (zepdrix):

Now just plug in your a values!

OpenStudy (anonymous):

\[a_3=\frac{a_1}{3}+\quad a_4=\frac{a_1}{4}+\quad a_5=\frac{a_1}{5}+\quad a_6=\frac{a_1}{6}+\quad a_7=\frac{a_1}{7}+\quad a_8=\frac{a_1}{8}+\quad a_9=\frac{a_1}{9}+\quad a_{10}=\frac{a_1}{10}+\quad a_{11}=\frac{a_1}{11}+\quad a_{12}\frac{a_1}{12}\] I'm sorry I'm really tired.... brain freeze...what do you mean by 'plug in your a values'

zepdrix (zepdrix):

What I posted a moment ago, is the FORM your solution will be in. \[a_1x^1+a_2x^2+a_3x^3+a_4x^4+...\]Plugging in the value for a_2 gives us,\[a_1x^1+\frac{a_1}{2}x^2+a_3x^3+a_4x^4+...\]

OpenStudy (anonymous):

sry :( rly tired oh my.... ok, so.... what is our goal?....I lost sight of it again.... Oh yeah... we're trying to write a sum describing this above pattern... We'll reach that goal by....

OpenStudy (anonymous):

I know I have some grammatical errors too...LOL

zepdrix (zepdrix):

lol :D poor gal

OpenStudy (anonymous):

\[y\quad =\quad \sum_{n=1}^\infty a_n x^n \quad =\] \[a_1x^1+a_2x^2+a_3x^3+a_4x^4+...\] ohhhh I plug in the a values!!!! haha

zepdrix (zepdrix):

Remember back to those friendly linear homogeneous DE's, we knew the answer would be of the form \[\large y=c_1 e^{r_1x}+c_2e^{r_2x}\]Something like that... Well with power series, we know that our solution will be of this form,\[y\quad =\quad \sum_{n=1}^\infty a_n x^n\] We expand out this power series, and from there we want to rewrite it as a series with ONLY 1 A value, not a bunch of different A values.

OpenStudy (anonymous):

\[y\quad =\quad \sum_{n=1}^\infty a_n x^n \quad =\] \[a_1x^1+\frac{a_1}{2}x^2+\frac{a_1}{3}x^3+\frac{a_1}{4}x^4+\frac{a_1}{5}x^5+\frac{a_1}{6}x^6+\frac{a_1}{7}x^7+\frac{a_1}{8}x^8+\frac{a_1}{9}x^9+\frac{a_1}{10}x^{10}\]

zepdrix (zepdrix):

Yah looks good c: So every term is written in terms of A1. So now we can rewrite it as a nice clean summation.

zepdrix (zepdrix):

Factoring out the A1 before we do that though*

OpenStudy (anonymous):

\[a_1\left(x^1+\frac{x^2}{2}+\frac{x^3}{3}+\frac{x^4}{4}+\frac{x^5}{5}+\frac{x^6}{6}+\frac{x^7}{7}+\frac{x^8}{8}+\frac{x^9}{9}+\frac{x^{10}}{10}\right)\]

zepdrix (zepdrix):

Notice a pattern? :) It's a nice simple one this time.

zepdrix (zepdrix):

\[\huge =a_1 \sum_{n=1}^\infty \frac{x^?}{?}\]

OpenStudy (anonymous):

\[y\quad =\quad a_1\sum_{n=1}^\infty \frac{x^n}{n} \quad \]

zepdrix (zepdrix):

Yayyyyy \:D/

zepdrix (zepdrix):

Hopefully we didn't make any mistakes along the way :U You never know!

OpenStudy (anonymous):

we forgot a negative sign according to chegg.com

OpenStudy (anonymous):

nope...sorry....my fault...we're right

zepdrix (zepdrix):

We are? :o oh yay.

zepdrix (zepdrix):

Oh ok, so we ARE suppose to start at 0, and then we just leave the A0 alone in front. I guess that makes sense :)

OpenStudy (anonymous):

I'll revisit that in the morning thanks for all of your help!!!

zepdrix (zepdrix):

np \c:/ haf a good nite

zepdrix (zepdrix):

and Merry Christmas! :D

OpenStudy (anonymous):

\[\huge \text{YOU} \quad \text{TOO}\quad =D\]